OpenStudy (anonymous):

Can anyone please help?? I'm not sure how to find the slope and y intercept of the line specified by the equation which is 2x+3-9=0?

OpenStudy (blurbendy):

you can find two points pretty easily. if you solve for x, you'll find that you end up with x = 3 now, replace 0 with y, and set x = 0 y = 2(0) + 3 -9 y = -6, when x = 0 so you have two points: (3, 0), (0, -6) slope = y2 - y1 / x2 - x1 -6 - 0 / 0 - 3 = -6 / -3 slope = 2

OpenStudy (blurbendy):

it might help if you draw the graph and plot the points to visualize everything